Circular No. 9068

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 Q2 (LINEAR-NEAT) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

comet P/2003 XD_{10} (cf. IAUC 8257) showing stellar appearance on Spacewatch images obtained with the 1.8-m f/2.7 reflector at Kitt Peak (first observation tabulated below).

2009 UT α_{2000} δ_{2000} Mag. Observer
Aug. 27.39043 1^{h}18^{m}00.^{s}41 −7^{o}20′05′′.0 21.5 Scotti

The indicated correction to the prediction on MPC 59598 is ΔT = −0.10 day. Additional astrometry by Scotti for Aug. 27, 28, and 29 — together with orbital elements for epoch 2003 (and those below), residuals, and an ephemeris — appear on MPEC 2009-Q68.

Epoch = 2010 Feb. 13.0 TT

T = 2010 Jan. 31.8363 TT  ω = 16.^{o}0938
e = 0.416377  Ω = 40.5285 }2000.0
q = 1.989593 AU  i = 13.4361

a = 3.409038 AU n^{o} = 0.1565872 P = 6.29 years

COMET C/2009 Q3 (LINEAR) J. V. Scotti, Lunar and Planetary Laboratory, reports his recovery of

comet P/2002 T1 (cf. IAUC 7983) showing a “soft” appearance (no obvious coma or tail) on Spacewatch images obtained with the 1.8-m f/2.7 reflector at Kitt Peak (first observation tabulated below).

2009 UT α_{2000} δ_{2000} Mag. Observer
Aug. 28.49035 4^{h}31^{m}39.^{s}80 +25^{o}09′34′′.5 21.1 Scotti

The indicated correction to the prediction on MPC 56804 is ΔT = +0.04 day. Additional astrometry by Scotti for Aug. 28 and 29 — together with orbital elements for epoch 2002 (and those below), residuals, and an ephemeris — appear on MPEC 2009-Q69.

Epoch = 2009 Sept. 6.0 TT

T = 2009 Aug. 25.5302 TT  ω = 3.^{o}8270
e = 0.639260  Ω = 14.2249 }2000.0
q = 1.314740 AU  i = 21.3965

a = 3.644561 AU n^{o} = 0.1416564 P = 6.96 years

2009 August 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green